# Assessing High-Risk Systems: An EU AI Act Verification Framework

### v1

@@ Abstract.
A central challenge in implementing the AI Act and other AI-relevant regulations in the EU is the lack of a systematic approach to verify their legal mandates. Recent surveys show that this regulatory ambiguity is perceived as a significant burden, leading to inconsistent readiness across Member States. This paper proposes a comprehensive framework designed to help close this gap by organising compliance verification along two fundamental dimensions: the type of method (controls vs. testing) and the target of assessment (data, model, processes, and final product). Additionally, our framework maps core legal requirements to concrete verification activities, serving as a vital bridge between policymakers and practitioners, and aligning legal text with technical standards and best practices. The proposed approach aims to reduce interpretive uncertainty, promote consistency in assessment practices, and support the alignment of regulatory, ethical, and technical perspectives across the AI lifecycle.
@@ 1. Introduction
The rapid advancement of Artificial Intelligence (AI) has created significant opportunities alongside new technical, organisational, and regulatory challenges, intensifying the need for trustworthy and well-governed systems. Recent regulatory initiatives, most notably the EU AI Act (eu_ai_act_2024), introduce a comprehensive set of legal obligations for high-risk AI systems. However, a persistent gap remains between these normative requirements and the technical and organisational means available to demonstrate and verify compliance in practice. Translating high-level principles such as fairness, robustness, and transparency into concrete, measurable testing and assurance activities continues to be challenging, requiring structured methodologies, interoperable assessment frameworks, and sustained collaboration across technical, legal, and ethical domains.
In parallel, uncertainty surrounding the Act’s practical implementation is increasingly recognised as a major challenge. This uncertainty manifests at multiple levels: interpretive uncertainty regarding the meaning and scope of specific legal obligations; operational uncertainty concerning their application in everyday development, deployment, and risk management practices; and procedural uncertainty linked to evolving standards, guidance documents, and the gradual designation of notified bodies (deloitte2024ai; dlapiper2025pause; draghi2024competitiveness; arnal2024implementation; lewis2025regulatorylearning; regulatinguncertainty2025). Together, these layers of uncertainty highlight the need for operational tools capable of translating legal expectations into consistently verifiable technical and procedural activities.
The motivation of this paper is to help bridge this gap between regulators, risk managers, developers, technical testers, and certifiers, who often operate with different vocabularies, processes, and levels of maturity. Establishing a shared operational framework can facilitate communication, reduce duplication of effort, and support coordinated assurance practices across the AI lifecycle.
The paper focuses on high-risk AI systems, which are subject to the most extensive obligations under the Act. Prohibited practices are banned outright, while low-risk systems are primarily subject to limited transparency duties that are largely subsumed within the requirements for high-risk systems. Concentrating on this category allows for a detailed examination of how legal obligations can be decomposed into testable, verifiable, and documentable elements, forming the basis for scalable and auditable compliance approaches.
Against this background, the paper is structured around three guiding questions: how high-level legal obligations can be systematically translated into operational components that can be tested, verified, or documented; which dimensions, artefacts, and assessment methods are needed to support a shared operational language among regulators, developers, auditors, and deployers; and how a unified assessment structure can enhance comparability, traceability, and communication across heterogeneous stakeholders and use cases.
The framework proposed in this paper does not aim to be complete or definitive. Rather, it offers a structured methodology intended to evolve alongside regulatory guidance, standardisation efforts, and advances in evaluation techniques. Its purpose is to provide a coherent starting point that can be progressively refined as institutional capacities, testing infrastructures, and sector-specific practices mature.
In addition to presenting the framework at a conceptual level, the paper demonstrates how it operates in practice through a concrete, real-world use case. By applying the framework to an ongoing high-risk AI system in the automotive domain, the paper illustrates how legal obligations can be mapped to technical and procedural controls across the system lifecycle. Overall, the paper presents a structured mapping of the Act’s obligations to currently available technical and organisational mechanisms, organised along four dimensions—data, model, processes, and final products—each associated with specific controls and testing activities.
2. Background
The AI Act introduces a risk-based framework with four categories: prohibited practices, high-risk systems, limited-risk systems, and minimal-risk systems. High-risk systems, defined in Article 6 and Annex III, concern domains where failures could affect health, safety, or fundamental rights, including critical infrastructure, employment, law enforcement, and justice administration. Providers must comply with extensive lifecycle obligations covering risk management, data quality, documentation, transparency, human oversight, and requirements for accuracy, robustness, and cybersecurity, with compliance demonstrated through internal or third-party conformity assessments depending on the system. Translating these obligations into verifiable technical criteria remains challenging. Concepts such as appropriate accuracy, sufficient robustness, or reasonably foreseeable misuse require contextual interpretation grounded in specific use cases and risk profiles. Dataset quality involves dimensions such as representativeness, label accuracy, and temporal stability, each implying different validation tasks. Robustness spans adversarial resilience, distributional stability, and degradation behaviour, necessitating distinct testing methods. Moreover, the Act’s lifecycle logic requires continuous monitoring, not merely point-in-time certification.
Although the technical community has developed numerous fairness, robustness, and transparency tests, they remain fragmented and only loosely connected to regulatory obligations. Recent initiatives provide partial structure: NIST’s AI RMF (NIST_AI_RMF_2025) addresses governance and measurement; ISO/IEC and IEEE standards offer terminology and risk-management guidance (iso_iec_42001_2023); and CEN-CENELEC JTC 21 (CEN-CENELEC_JTC21_2025) is developing harmonized European standards that, under Article 40, will grant presumption of conformity. This work is ongoing, creating a transitional period in which interim frameworks are needed to orient compliance. Notified bodies for third-party assessments (Article 34) are in the process of being designated, while regulatory sandboxes (Article 57) offer controlled environments to test innovative systems and refine assessment approaches.
Persistent challenges remain: misaligned terminology across legal, ethical, and technical communities; fragmented testing methods lacking integration; limited procedural guidance for assessments; insufficient attention to post-deployment phases; and the absence of standardised approaches enabling comparability across systems. The framework proposed here addresses these gaps by offering a structured bridge between the Act’s normative requirements and practical validation processes (mowla2024ai).
@@ 3. Methodology
This section outlines the methodological process used to construct the proposed framework, detailing how requirements were identified, categorized, and systematically linked to regulatory, standardization, and scientific sources. The aim is to ensure that every control and testing mechanism rests on a coherent chain of traceability, from high-level legal principles to concrete, technically verifiable methods.
3.1. Requirements
We identified 11 macro-categories of requirements that collectively capture the key dimensions of AI trustworthiness and compliance. The starting point for this categorization was the seven principles of Trustworthy AI defined by the European Commission’s High-Level Expert Group on Artificial Intelligence (HLEG) in its Ethics Guidelines for Trustworthy AI (eu_trustworthy_ai_2019). These principles, formulated prior to the adoption of the EU AI Act, have long served as a foundational reference for aligning AI development with European ethical and societal values. Although the AI Act introduces a more prescriptive and legally binding framework, the essence of these principles remains embedded throughout the regulation: in some cases as stand-alone articles (e.g., transparency, human oversight), and in others distributed across multiple provisions concerning governance, documentation, and system performance.
To complement these ethical foundations with more operational and procedural dimensions, we added four additional categories corresponding to key obligations introduced by the AI Act: quality management, risk management, technical documentation, and record keeping. These categories address the infrastructural and organizational mechanisms that ensure traceability, consistency, and accountability across the AI lifecycle. Together, the original HLEG principles and these four operational domains constitute the 11 macro-requirements underpinning the proposed framework.
Building upon this foundation, each macro-category was further refined into a set of more granular requirements. These were derived through a close reading of the AI Act’s legal provisions, complemented where necessary by interpretative inference from the broader normative intent of the macro-level requirements. This two-tiered decomposition, from principles to operational requirements, enabled a systematic and traceable linkage between high-level ethical concepts and their concrete technical or procedural manifestations.
Once the granular requirements were defined, we proceeded to map each of them to specific guidelines and methods. This mapping was performed hierarchically: first connecting each requirement to relevant guidance sources, and subsequently identifying testing or control methods capable of empirically or procedurally validating compliance. This structured mapping approach ensures conceptual continuity across levels, from principle to requirement, from requirement to guideline, and from guideline to method, thereby providing a coherent foundation for both qualitative governance assessment and quantitative system testing.
3.2. Sources
The identification of control and testing methods followed a structured, multi-stage approach aimed at ensuring both regulatory relevance and technical completeness. The objective was to derive mechanisms that are not only aligned with current legal and ethical requirements, but also grounded in established standards and empirically validated research.
First, we conducted a systematic examination of binding and quasi-binding regulatory sources, beginning with the EU AI Act (eu_ai_act_2024) and the General Data Protection Regulation (GDPR)  (GDPR2016). These texts define the normative baseline for high-risk AI systems in the European context, providing explicit legal obligations related to transparency, robustness, and accountability. Their provisions were used to delimit the scope of each requirement and to establish traceability between legal articles and technical mechanisms.
Second, we reviewed international standards and authoritative guidance documents issued by globally recognised bodies, such as the EU’s General Purpose AI Code of Practice  (gpai-code-practice-3rd), ISACA’s Advanced Audit in AI Official Review Manual (isaca2025aaia), UNESCO’s Recommendation on the Ethics of Artificial Intelligence (unesco2021ethics), ISO/IEC 42001:2023 (iso_iec_42001_2023), ISO23894 (ISO23894) and ISO9001 (ISO9001). These documents were selected because they operationalize high-level principles into implementable governance or audit processes and therefore provide an essential bridge between normative and technical domains.
Third, where specific mappings between legal principles and technical procedures were not sufficiently covered by regulatory or standardization sources, we extended the analysis to the scientific literature. We focused on highly recognised and widely cited works in AI governance, robustness testing, and risk management (e.g., (Brundage2020; Basiri2016; Floridi2018; mundhenk2020)). Inclusion criteria combined citation count, influence of the publication venue, and the relevance of the paper to measurable aspects of trustworthiness such as fairness, reliability, or transparency. This ensured that the resulting mappings draw from the most authoritative and empirically supported methodologies currently available.
Through this layered approach, progressing from regulations to standards to scientific evidence, we aimed to guarantee both the normative validity and the technical soundness of the proposed framework. Each subsection that follows in Section 4 corresponds to one Trustworthy AI requirement and enumerates its associated control and testing mechanisms accordingly.
3.3. Dimensions of Assessment
Figure 1. Mapping between the assessment type and the assessment target dimensions.
Before mapping requirements to methods in Section 4, it is necessary to introduce the conceptual structure underpinning the framework. The analysis of an AI system can be meaningfully organised along two complementary dimensions of assessment. The first dimension concerns the type of assessment, The second dimension concerns the target of assessment. Together, these two axes define the space in which assurance, verification, and risk mitigation take place.
Along the type of assessment axis, the framework distinguishes between controls and testing. Controls are primarily ex-ante activities that establish governance, documentation, and procedural safeguards to ensure responsible development before deployment. Testing, by contrast, consists of ex-post evaluations that empirically examine system behaviour, performance, or statistical properties under specified conditions. Neither approach is hierarchical; rather, they are mutually reinforcing. Controls ensure that development processes are auditable and accountable, while testing provides empirical evidence that outcomes align with intended requirements.
Along the target of assessment axis, the framework identifies four lifecycle layers. The data layer captures aspects related to data governance such as the acquisition, composition, and representativeness of data. Model related concerns such as its architecture, explainability, robustness, and performance are covered by the model dimension. The processes layer encompasses organisational and procedural elements of system development and maintenance, including risk and quality management, documentation, traceability, and lifecycle governance. It reflects how the system is produced and managed rather than what it does. Finally, the final product layer focuses on the behaviour of the deployed system, its interaction with users, and its broader societal or environmental implications, where outputs and impacts can be empirically observed and validated.
The intersection of these two axes yields a comprehensive assessment model that connects organisational assurance with empirical verification across all lifecycle stages. This structure enables evaluators to determine, for each obligation under the AI Act, which combination of assessment type and target is most appropriate, ensuring that governance and verification are treated as interdependent components of a coherent assurance continuum.
@@ 4. Mapping
This section presents the mapping between the macro-requirements of Trustworthy AI and the corresponding mechanisms that enable their implementation and verification. Each subsection (R1–R11) details one requirement category, identifying the relevant legal provisions, associated guidelines, and the methods through which compliance can be assessed or demonstrated.
To ensure conceptual clarity, each method is classified along the two analytical dimensions described in 3. Type distinguishes between C (Controls) and T (Tests). Target specifies the primary layer of the AI lifecycle to which the method applies: D (Data), M (Models) , P (Processes) , and FP (Final Product).

### v2

# Assessing High-Risk AI Systems under the EU AI Act: From Legal Requirements to Technical Verification

@@ Abstract.
The implementation of the AI Act requires practical mechanisms to verify compliance with legal obligations, yet concrete and operational mappings from high-level requirements to verifiable assessment activities remain limited, contributing to uneven readiness across Member States. This paper presents a structured mapping that translates high-level AI Act requirements into concrete, implementable verification activities applicable across the AI lifecycle. The mapping is derived through a systematic process in which legal requirements are decomposed into operational sub-requirements and grounded in authoritative standards and recognised practices. From this basis, verification activities are identified and characterised along two dimensions: the type of verification performed and the lifecycle target to which it applies. By making explicit the link between regulatory intent and technical and organisational assurance practices, the proposed mapping reduces interpretive uncertainty and provides a reusable reference for consistent, technology-agnostic compliance verification under the AI Act.
@@ 1. Introduction
Progress in AI has created substantial opportunities while also introducing significant technical, organisational, and regulatory challenges, particularly as AI systems are increasingly deployed in critical domains where failures may have serious societal consequences (bengio2024managing; wang2022artificial). In response, regulatory efforts worldwide are intensifying to promote trustworthiness in AI development and deployment (huang2024; smuha2021race; novelli2024taking). Within this context, the EU Artificial Intelligence Act (AI Act) (eu_ai_act_2024) represents one of the most comprehensive regulatory responses to date, establishing a risk-based framework that mandates ex ante verification and ongoing oversight for high-risk AI systems. Despite this progress, a persistent gap remains between high-level legal requirements and the concrete verification activities needed to demonstrate compliance in practice. This gap reflects the absence of explicit and shared mappings between regulatory obligations and implementable assurance practices, resulting in fragmented compliance efforts across stakeholders and Member States and limiting comparability and regulatory learning. In practice, organisations often struggle to translate abstract legal obligations into verifiable controls, tests, and documentation artefacts that can be applied consistently across the AI lifecycle.
These challenges are compounded by three interrelated forms of uncertainty. Interpretive uncertainty concerns the contextual meaning and scope of legal obligations; operational uncertainty relates to how such obligations can be instantiated across heterogeneous development and deployment practices; and procedural uncertainty arises from evolving standards, guidance documents, and the gradual designation of notified bodies (deloitte2024ai; dlapiper2025pause; draghi2024competitiveness; arnal2024implementation; lewis2025regulatorylearning; regulatinguncertainty2025). These uncertainties are further amplified by the emergent behaviour of modern generative AI systems, whose properties arise from complex interactions rather than explicit programming (holtzmanGenerativeModelsComplex2025; koltLessonsComplexSystems2025a). As a horizontal regulation relying on harmonised standards, post-market monitoring, and iterative guidance, the AI Act embodies a model of regulatory learning in which compliance expectations are progressively refined through implementation experience and verification outcomes (zakiConceptualisingOrganisationalPolicy2025; lewis2025regulatorylearning; schrepelAdaptiveRegulation2025). Such learning depends on the availability of structured, comparable, and traceable verification evidence, which remains fragmented across actors, disciplines, and use cases.
This paper addresses the need for an explicit and operational mapping between AI Act requirements and concrete verification activities that can be executed, documented, and compared across the AI lifecycle. The objective is to operationalise compliance by making the link between regulatory intent and technical and organisational assurance practices explicit, reproducible, and auditable. The focus is on high-risk AI systems as defined in Article 6 and Annex III, which concern application domains where failures may adversely affect health, safety, or fundamental rights. Compliance must be demonstrated through conformity verification procedures under Articles 43 and 44, either via internal control procedures (Annex VI) or third-party assessment by notified bodies (Annex VII), both requiring verifiable evidence spanning design, development, deployment, and post-market monitoring.
Translating these obligations into concrete assessment activities remains challenging. Many provisions of the AI Act are inherently context-dependent, requiring case-specific interpretation grounded in the system’s intended use, operational environment, and evolving risk profile. Requirements related to data governance, robustness, accuracy etc. encompass multiple dimensions, each necessitating distinct validation and testing methodologies, while the lifecycle-oriented nature of the Act entails continuous monitoring and post-deployment evaluation rather than reliance on point-in-time certification alone. Existing initiatives provide partial foundations for AI verification, but none offers a complete and operational mapping between legal requirements and verification activities. The NIST AI Risk Management Framework (NIST_AI_RMF_2025) provides high-level guidance but is not aligned with EU-specific legal obligations, while ISO/IEC standards such as ISO/IEC 42001 (iso_iec_42001_2023) establish process-oriented management requirements without specifying AI Act–specific verification procedures. European standardisation efforts within CEN–CENELEC JTC 21 (CEN-CENELEC_JTC21_2025) aim to develop harmonised standards granting presumption of conformity under Article 40, but this work remains ongoing.
Against this background, the central contribution of this paper is a structured and operational mapping that translates high-level AI Act requirements into concrete verification activities applicable across the AI lifecycle. The mapping is constructed through a systematic decomposition of legal obligations into operational sub-requirements and is grounded in authoritative standards and recognised engineering and governance practices. To ensure consistency and comparability, verification activities are characterised along two dimensions capturing both verification type and lifecycle target. The applicability of the mapping is illustrated through a real-world case study involving a high-risk AI system in the automotive domain. Within this scope, the paper addresses three guiding questions that structure the construction of the proposed mapping: (i) how high-level legal obligations can be decomposed into concrete verification activities; (ii) which verification dimensions are required to ensure comparability and traceability across heterogeneous actors; and (iii) how a shared verification structure can support communication and regulatory learning across use cases and sectors. While related efforts exist (e.g. (hernandez2025open)), the proposed mapping provides a higher level of granularity and operationalisability. It is not intended to be prescriptive or exhaustive, but to serve as a reusable reference that can evolve alongside regulatory guidance, standardisation efforts, and advances in evaluation techniques, thereby supporting more consistent conformity verification and evidence-based regulatory learning over time.
@@ 2. Methodology
This section describes the methodology used to construct a structured mapping between high-level AI Act requirements and concrete, implementable verification activities. Figure 1 provides an overview of the resulting compliance verification structure and the main methodological building blocks. It illustrates how high-level legal requirements are progressively decomposed, grounded, and translated into concrete verification activities, which are then situated within a structured verification space defined along two core dimensions.
Figure 1. Overview of the methodology.
2.1. Normative inputs and requirement structuring
The starting point of the methodology consists in identifying and structuring a set of high-level requirements derived from the obligations established by the AI Act. Given the heterogeneity and dispersion of these obligations across the regulation, an explicit structuring step is required to organise them into a coherent form suitable for systematic verification. For this purpose, we use the seven principles of Trustworthy AI defined by the European Commission’s High-Level Expert Group in the Ethics Guidelines for Trustworthy AI (eu_trustworthy_ai_2019) as an organising abstraction. Although these principles predate the AI Act and are not legally binding, they have strongly influenced the development of European AI governance and are reflected throughout the regulation, both as explicit obligations (e.g. transparency, human oversight) and as distributed requirements embedded in provisions on governance, documentation, and system performance. Their role in the methodology is therefore structural: they serve as a stable taxonomy for grouping and analysing legally binding requirements, not as an independent source of obligations.
To account for the more prescriptive and organisational character of the AI Act, this abstraction is complemented with four additional requirement categories corresponding to obligations that are central to conformity assessment for high-risk systems: quality management, risk management, technical documentation, and record keeping. These categories capture lifecycle governance and accountability mechanisms that are explicitly mandated by the regulation but are only partially represented in principle-based formulations. Together, the resulting eleven high-level requirements provide a structured normative representation of the obligations imposed by the AI Act for high-risk systems. Each requirement corresponds to a coherent cluster of legally binding provisions, organised in a form suitable for systematic verification and without introducing additional normative content beyond the regulation itself.
Each high-level requirement is subsequently decomposed into more granular operational components. These components correspond either to obligations explicitly stated in the AI Act or to requirements that can be reasonably derived from its provisions in order to enable verification. This decomposition does not aim to provide an exhaustive or definitive interpretation of the law, but to articulate traceable operational interpretations that can be subjected to assessment. The methodology explicitly allows for alternative decompositions depending on context, sector, and risk profile, and is designed to accommodate such variability rather than eliminate it.
2.2. Authoritative grounding
To ensure both normative alignment and technical soundness, the operational interpretation of requirements is grounded in authoritative sources that specify or exemplify how high-level obligations are implemented and assessed in practice. These sources are predominantly non-binding, complemented where appropriate by binding instruments whose established operational mechanisms are widely reused in compliance and assurance activities. In this respect, the GDPR (GDPR2016) is included not as a source of normative extension of the AI Act, but because it provides mature and widely operationalised mechanisms—such as risk-based assessment, documentation practices, and accountability procedures—that directly inform how several AI Act obligations are interpreted and verified in practice. The AI Act explicitly operates in complementarity with existing data protection law, and high-risk AI systems in the European context frequently involve personal data processing throughout development, deployment, and monitoring. As a result, verification activities related to data governance, transparency and information provision, record keeping, and affected-person rights often rely on GDPR-derived concepts, artefacts, and procedures, such as information duties, accountability documentation, and impact-assessment practices. For this reason, the GDPR is treated here as an authoritative operational reference within the grounding layer.
Beyond this exception, operational grounding relies on non-binding but authoritative sources that provide reusable mechanisms for implementing and assessing regulatory obligations in practice. These include the EU General Purpose AI Code of Practice (gpai-code-practice-3rd), ISACA’s Advanced Audit in AI Official Review Manual (isaca2025aaia), UNESCO’s Recommendation on the Ethics of AI (unesco2021ethics), the Web Content Accessibility Guidelines (WCAG) 2.2 (wcag22), and a broad range of ISO and IEC standards relevant to AI governance, risk management, quality management, cybersecurity, and data quality (e.g. (ISO_IEC_22989_2022; ISO23894; ISO_IEC_TR_24027_2021; ISO_IEC_24029_2_2023; ISO_IEC_24970_DIS_2025; ISO_IEC_27001_2022; ISO_IEC_27005_2022; ISO_IEC_27035_1_2023; ISO9001; ISO_31000_2018; ISO_IEC_31010_2019; iso_iec_42001_2023; ISO_IEC_5259_1_2024; ISO_8000_8_2015)). These sources were selected because they translate high-level principles into implementable governance processes, controls, and audit practices, thereby bridging normative requirements and operational verification.
Where regulatory, guidance, and standardisation sources did not sufficiently specify technical verification procedures, the analysis was extended to the scientific literature. Highly cited and influential works in AI governance, robustness testing, risk management, and human–system interaction were considered (e.g. (Brundage2020; Basiri2016; Floridi2018; mundhenk2020; sheridan1992telerobotics; schuler1993participatory; CoE_AI_2020)). Selection criteria combined citation impact, venue relevance, and explicit connection to measurable or testable aspects of AI trustworthiness. This layered grounding process ensures that each operational requirement and associated verification activity can be traced back to authoritative legal, standardisation, or scientific sources.
2.3. Verification dimensions
Figure 2. Verification space defined by verification type and verification target.
Concrete verification activities are organised within a structured verification space defined by two orthogonal dimensions: the type of verification and the target of verification within the AI lifecycle. Together, these dimensions provide a common operational language for identifying, organising, and comparing verification activities across heterogeneous systems and stakeholders. The first dimension concerns the type of verification. Two broad categories are distinguished: controls and testing. Controls refer to process-based assurance mechanisms, including governance structures, documentation practices, quality management procedures, and organisational safeguards that ensure responsible system development and operation. Testing refers to empirical evaluations that assess system behaviour, performance, or properties under specified conditions. Controls and testing are complementary rather than hierarchical: controls establish accountability and traceability, while testing provides empirical evidence of compliance. Both may be applied at any stage of the AI lifecycle. The second dimension concerns the target of verification, corresponding to the lifecycle component or artefact being evaluated. Numerous software and AI lifecycle models provide fine-grained decompositions of development and deployment phases (e.g. (de2022artificial; wang2021artificial)). However, such models vary significantly across standards, engineering practices, and organisational contexts, and their level of detail often presupposes technical expertise. For the purpose of compliance verification, which involves regulators, auditors, legal experts, risk managers, and developers alike, a higher level of abstraction is required to support immediate interpretability and shared understanding.
Accordingly, the methodology adopts four high-level verification targets that act as stable and intuitive aggregation buckets rather than exhaustive lifecycle stages. The data target addresses issues related to data acquisition, composition, representativeness, and quality. The model target concerns the AI system’s architecture, training procedures, explainability, robustness, and performance characteristics. The processes target encompasses organisational and procedural elements such as risk management, quality assurance, documentation, traceability, and lifecycle governance. The final product target focuses on the behaviour of the deployed system, including its interaction with users, integration into sociotechnical contexts, and observable outputs and impacts. This abstraction is intentional: it sacrifices fine-grained lifecycle specificity in favour of clarity, stability, and cross-disciplinary usability. Each target is sufficiently broad to accommodate different development methodologies and technical implementations, while remaining concrete enough to support systematic verification and evidence collection.
While analytically distinct, these targets may overlap in practice. Certain verification activities, such as model monitoring or data versioning, span multiple targets simultaneously. The methodology treats these boundaries as pragmatic rather than absolute, allowing verification activities to address multiple concerns where appropriate. Within this verification space, each high-level requirement is associated with a set of control and testing activities consistent with its legal intent and operational scope. This structure enables evaluators to systematically identify relevant verification activities, organise evidence coherently, and ensure coverage across governance and empirical dimensions. By making explicit how regulatory requirements translate into concrete verification actions, the methodology supports comparability, traceability, and communication across heterogeneous stakeholders and use cases.
@@ 3. Mapping
In this section, we present the mapping between the high-level requirements of the AI Act defined in Section 2.1 and the mechanisms that support their implementation and verification. Each subsection (R1–R11) introduces a brief explanation of the corresponding requirement, followed by the actual mapping presented in the form of a table. These tables identify the relevant legal provisions, associated authoritative references, and the concrete methods through which compliance can be assessed or demonstrated. To ensure conceptual clarity and consistency across requirements, each method is classified along the two analytical dimensions described in Section 2.3. Type distinguishes between C (Controls) and T (Tests), while Target specifies the primary layer of the AI lifecycle to which the method applies: D (Data), M (Models), P (Processes), and FP (Final Product).